OpenStudy (anonymous):

The legs of a right triangle have lengths of 4 and 5 units. Find the length of the hypotenuse. 3 3sqrt3 sqrt 41

Nnesha (nnesha):

 hypotenuse is the longest side of right triangle \[\huge\rm a^2+b^2= c^2\] c is hypotenuse (longest side) replace a and by their value

Nnesha (nnesha):

btw 4 ,5,?, is special triangle :-)

OpenStudy (anonymous):

9^2?

Nnesha (nnesha):

how did you get 9 ?

OpenStudy (anonymous):

added 4+5

Nnesha (nnesha):

A^2 = ? B^2 = ? then add both numbers

OpenStudy (anonymous):

got it thnx

Nnesha (nnesha):

so what's ur answeR ?

OpenStudy (anonymous):

41

Nnesha (nnesha):

a can be 4 or 5 doesn't matter a=4 b=5 4^2 is same as 4 times 4 = ?? 5^2= ??

OpenStudy (anonymous):

are you tying to tell me that the answer isn't sqrt 41?

Nnesha (nnesha):

\[\huge\rm 4^2 +5^2 =c^2\] i'm trying to tell you how to solve this question

OpenStudy (anonymous):

ya I got it

Nnesha (nnesha):

we'll see at the end

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @hhelpplzzzz ya I got it \(\color{blue}{\text{End of Quote}}\) nope

Nnesha (nnesha):

1st)take square of both numbers 2nd) add 3rd) take square root